Tell whether the following sets A are (i) open or closed, (ii) connected,




(iii) compact. Find (iv) the limit points of A, (v) int(A), (vi) the boundary of A.





A= rational numbers




A = integers






1
Expert's answer
2022-02-08T16:31:16-0500

First let us start with the case "A= \mathbb{Q}\subset \mathbb{R}"

  1. "A" is neither open (as any neighbourhood of any point contains an irrational number) nor closed (as a limit of rational numbers need not to be rational).
  2. "A" is not connected, we can consider two following sets : "A^1 = (-\infty;\sqrt{2}), A^2 = (\sqrt2 ; +\infty)". Both of them are open and "A\subseteq A^1\cup A^2".
  3. "A" is not compact, as it is not closed in "\mathbb{R}" (any compact subset of "\mathbb R" is, in particular, a closed subset).
  4. Limit points of "A" is the entire real line "\mathbb{R}", as any real number is a limit of a sequence of rational numbers.
  5. "\text{int} A = \empty", as any non-empty open set contains an irrational number, and so no non-empty open set is contained in "A".
  6. "\partial A = \mathbb{R}", as the boundary is exactly the closure of "A" minus the interior of "A", i.e. "\mathbb{R} \setminus \emptyset = \mathbb R".

Now let us consider the case "A = \mathbb Z \subseteq \mathbb R"

  1. "A" is closed, as any converging sequence of integers should be stable, so it converges to an integer. "A" is not open, as any neighbourhood of an integer contains non-integer numbers.
  2. "A" is not connected, we can consider the sets "A^1=(-\infty; \frac{1}{2}), A^2=(\frac{1}{2}, +\infty)". Both of them are open and "A\subseteq A^1 \cup A^2".
  3. "A" is not compact, as the sequence "x_i = i" does not admit a converging subsequence. Alternatively, the open cover "(i-\frac{1}{2}, i+\frac{1}{2})" for "i\in \mathbb Z" does not admit a finite subcover.
  4. The limit points of "A" are empty, as no integer "n \in \mathbb Z" can be approximated by a sequence of other integers.
  5. "\text{int} A = \empty" as any non-empty open set contains a non-integer number.
  6. "\partial A = A = \mathbb{Z}", as "A" is a closed set with an empty interior.